Two complementary angles are in the ratio of 2:3, find the measure of the complementary angle.
2+3=5 90/5=18

Found 2 solutions by RAY100, Alan3354:
Answer by RAY100(1637)

Let one angle be 2x,,,and other angle be 3x
.
2x +3x = 90,,,,,,,def of complement
.
5x = 90
.
x=18
.
angles are 2x = 36, and 3x = 54
.
check,,,,,36 +54 = 90 ,,,,,,ok
.
36 /54 = 2/3,,,,,ok

Answer by Alan3354(69443)

Two complementary angles are in the ratio of 2:3, find the measure of the complementary angle.
2+3=5
90/5=18
-------
Then, 1 angle is 2*18, the other is 3*18
= 36º and 54º